# Resumming Post-Minkowskian and Post-Newtonian gravitational waveform expansions

### v2

@@ V.1 The hypergeometric representation
$G_{\alpha}(z)=G^{0}_{\alpha}({\textstyle\frac{x}{z}})g_{\alpha}(x)+\ldots=G^{1}_{\alpha}(z)+\ldots$ (92)
@@
$H^{0}_{\alpha}(y)$ $=$ $y^{\kappa_{1}-\alpha a}{}_{2}F_{1}(\kappa_{1}{-}\alpha a,1{+}\kappa_{1}{-}\gamma_{0}{-}\alpha a,1{-}2\alpha a,y)$
$H^{1}_{\alpha}(z)$ $=$ $z^{\alpha a-\kappa_{1}}{}_{2}F_{1}\left(\alpha a{+}\kappa_{2},\alpha a{+}\kappa_{3};1{+}2\alpha a;z\right)$ (93)
@@
$P_{0}(y)$ $=$ $1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\,,\,\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i}\widehat{c}_{ij}\,x^{i}y^{j-i}$
$P_{1}(z)$ $=$ $1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}\,,\,\widehat{P}_{1}(z)=\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}$
We notice that now only the non-confluent hypergeometric functions are involved. As before, plugging the ansätze (92) into (86) and setting to zero the coefficients of $H^{p}_{\alpha}$ ($p=0,1$) and its derivatives, one finds a linear system of algebraic equations for the coefficients $c_{ij}$, $\widehat{c}_{ij}$, $d_{ij}$, $\widehat{d}_{ij}$. For example, for $k=1$ one finds
$c_{10}=$ $\frac{\left[2a^{2}\left(\kappa_{1}{+}\kappa_{2}{+}\kappa_{3}{-}\frac{1}{2}\right){+}2\kappa_{1}\kappa_{2}\kappa_{3}{-}\kappa_{1}\kappa_{2}{-}\kappa_{1}\kappa_{3}{-}\kappa_{2}\kappa_{3}\right]}{4a^{2}{-}1}$
$\widehat{c}_{11}=$ $-\widehat{c}_{10}=\frac{\left(2a^{2}-\kappa_{2}-\kappa_{3}+2\kappa_{2}\kappa_{3}\right)}{4a^{2}{-}1}\qquad\qquad\qquad$
$d_{01}=$ $1-\frac{y\left(2\gamma_{0}{-}2\kappa_{1}{-}1\right)\left(a^{2}{-}\kappa_{1}^{2}\right)}{4a^{2}-1}\qquad\qquad\qquad\quad$
$\widehat{d}_{11}=$ $-\widehat{d}_{01}=\frac{\left(2\kappa_{1}^{2}{+}2a^{2}{-}2\gamma_{0}\kappa_{1}{+}\gamma_{0}{-}1\right)}{4a^{2}{-}1}\qquad\qquad\quad$
and, writing $\widehat{u}=a^{2}+\sum_{i}\widehat{u}_{i}x^{i}$, we get
$\widehat{u}_{1}=\frac{\gamma_{0}\left[a^{2}\left(1-2\kappa_{1}-2\kappa_{2}-2\kappa_{3}\right)+\kappa_{1}\kappa_{2}+\kappa_{1}\kappa_{3}+\kappa_{2}\kappa_{3}-2\kappa_{1}\kappa_{2}\kappa_{3}\right]}{4a^{2}-1}-\frac{\left(a^{2}-\kappa_{1}^{2}\right)\left(2a^{2}-\kappa_{2}+2\kappa_{2}\kappa_{3}-\kappa_{3}\right)}{4a^{2}-1}\$ (95)
As in the confluent case, we denote by $g_{\alpha}$ the ratio $G^{1}_{\alpha}(z)\over G^{0}_{\alpha}({\textstyle\frac{x}{z}})$. Explicit computations show that, as in the confluent case, the ratio of the two components can be written as
with ${\cal F}_{\rm inst}$ still is given by (18), (12) and
$u(a)=\frac{\widehat{u}(a)+x\left[\kappa_{2}\kappa_{3}+(\gamma_{0}-\kappa_{1})(\kappa_{2}+\kappa_{3})\right]}{1-x}$ (97)
The function ${\cal F}_{\rm inst}$ corresponds now to the prepotential of a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with four fundamental flavours.
V.2 Connection formulae
@@
$\widetilde{H}^{0}_{\alpha}(y)$ $=$ $e^{{\rm i}\pi\frac{1{+}\alpha}{2}(1{-}\gamma_{0})}y^{\frac{1{+}\alpha}{2}(\gamma_{0}{-}1)}{}_{2}F_{1}\left({\textstyle\frac{1{+}\alpha}{2}}(1{-}\gamma_{0}){+}\kappa_{1}{+}a,{\textstyle\frac{1{+}\alpha}{2}}(1{-}\gamma_{0}){+}\kappa_{1}{-}a,1{+}\alpha\left(1{-}\gamma_{0}\right);y^{-1}\right)$
$\widetilde{H}^{1}_{\alpha}(z)$ $=$ $z^{a-\kappa_{1}}(1{-}z)^{{1{+}\alpha\over 2}(1{-}\kappa_{2}{-}\kappa_{3})}\,_{2}F_{1}\left({\textstyle\frac{1{+}\alpha}{2}}{+}a{-}\alpha\kappa_{2},{\textstyle\frac{1{+}\alpha}{2}}{+}a{-}\alpha\kappa_{3},1{+}\alpha\left(1{-}\kappa_{2}{-}\kappa_{3}\right);1{-}z\right)$ (98)
We notice that both $\widetilde{H}^{0}_{\alpha}(y)$ and $\widetilde{H}^{1}_{\alpha}(z)$ are invariant under the sign flipping $a\to-a$, due to hypergeometric identities. The functions (98) are related to those introduced in (93) by the standard hypergeometric relations
$H^{0}_{\alpha}(y)=\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{H}^{0}_{\alpha^{\prime}}(y),H^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{H}^{1}_{\alpha^{\prime}}(z)$ (99)
@@
$B_{\alpha\alpha^{\prime}}$ $=$ $\frac{e^{{\rm i}\pi(\kappa_{1}{-}\alpha a)}\Gamma(1{-}2\alpha a)\Gamma(\alpha^{\prime}(\gamma_{0}{-}1))}{\Gamma\left({\textstyle\frac{1{-}\alpha^{\prime}}{2}}{-}\alpha a{+}\alpha^{\prime}(\gamma_{0}-\kappa_{1})\right)\Gamma\left({\textstyle\frac{1{-}\alpha^{\prime}}{2}}{-}\alpha a{+}\alpha^{\prime}\kappa_{1}\right)}$
$F_{\alpha\alpha^{\prime}}$ $=$ $\frac{\Gamma(1{+}2\alpha a)\Gamma(\alpha^{\prime}[\kappa_{2}{+}\kappa_{3}{-}1])}{\Gamma\left({\textstyle\frac{1-\alpha^{\prime}}{2}}{+}\alpha a{+}\alpha^{\prime}\kappa_{2}\right)\Gamma\left({\textstyle\frac{1-\alpha^{\prime}}{2}}{+}\alpha a{+}\alpha^{\prime}\kappa_{3}\right)}$ (100)
Plugging (99) into (V.1) one finds the connection formulae
@@
$\widetilde{G}^{0}_{\alpha}(y)$ $=$ $\left[P_{0}(y)\widetilde{H}^{0}_{\alpha}(y){+}\widehat{P}_{0}({\textstyle\frac{x}{z}})y\widetilde{H}^{0}_{\alpha}{}^{\prime}(y)\right]$
$\widetilde{G}^{1}_{\alpha}(z)$ $=$ $\left[P_{1}(z)\widetilde{H}^{1}_{\alpha}(z){+}\widehat{P}_{1}(z)z\widetilde{H}^{1}_{\alpha}{}^{\prime}(z)\right]$ (102)
@@
VI Summary and conclusions
In this paper, we computed the waveform emitted by a particle moving in a Schwarzschild geometry under the assumption of low frequencies $x=4{\rm i}\omega M\ll 1$. There are two ways of taking this limit, depending on whether one keeps finite the variable $z=2M/r$ or $y=x/z=2{\rm i}\omega r$. We show that in both limits, the solution at order $x^{k}$ of the confluent Heun equation governing the radial motion can be written as the linear sum of a single hypergeometric (or confluent hypergeometric) function and its first derivative with coefficients given by polynomials of degree $k$, determined in a recursive fashion. The results provide an (all order in velocity) formula for the PM expansion, and an (all order in the coupling) formula for the PN series.
The parameters of the hypergeometric function and the coefficients of the two polynomials are all expressed in terms of a characteristic function $a(\ell)$, that plays the role of the SW period in a gauge theory description or the ”renormalized angular momentum” in the MST formulation. The connection formulae are derived from the standard hypergeometric transformation rules. The complexity of the Heun transformation rule is hidden in the non trivial dependence $a(\ell)$ (57) on the frequency and the angular momentum $\ell$.
The PM and PN representations $G^{0}_{\alpha}(z)$ and $G^{1}_{\alpha}(z)$ cover the exterior $r\gg 2M$ and interior $\omega r\ll 1$ regions of the spacetime. Given in terms of hypergeometric functions, they can be extrapolated all the way to radial infinity or to the horizon using the standard transformation rules for the hypergeometric functions. We denote by $\widetilde{G}^{p}_{\alpha}$ the bases of solutions describing these limits. Finally, we show that $G^{0}_{\alpha}(z)$ is proportional to $G^{1}_{\alpha}(z)$ and determine the coefficient of proportionality $g_{\alpha}$ order by order in $x$ by matching the two functions in the overlapping near zone. The final picture is summarized in the following scheme
@@ VI Summary and conclusions
$\begin{array}[]{lll}{\rm Near\leavevmode\nobreak\ horizon}&{\rm Near\leavevmode\nobreak\ zone}&{\rm Far\leavevmode\nobreak\ zone}\\ z\approx 1&x\ll z\ll 1&z\ll x\ll 1\\ F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}&G^{1}_{\alpha}=g_{\alpha}G^{0}_{\alpha}&g_{\alpha}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}\end{array}$ (107)
@@
with $B_{\alpha\alpha^{\prime}}$ and $F_{\alpha\alpha^{\prime}}$ the hypergeometric connection matrices sending $z\to 1/z$ and $z\to z+1$ respectively. The whole non-triviality of the Heun connection formula is codified in the dependence of these matrices on the characteristic function $a(\ell)$.
@@
The incoming and upgoing waveforms are given by linear combinations of any of these bases, see for example (69). When compared against the PN expansions that one can find in literature [7, 20], these formulae resum the infinite tower of logarithmically divergent terms (tail and tail of tails), responsible for the slow convergence of the PN series, into exponentials. They also summarise in a compact form the dependence on $\ell$ of the waveforms, all the so called tail and tail of tail terms of the multipole expansion and have no logarithmic terms (responsible for the slow convergence of the PN series). The results for the waveform have been compared against those obtained via MST and numerical methods, finding an agreement from 9 to 13 digits in the whole range of parameters allowed for bounded orbits.
@@
Similar results are obtained for solutions of the non-confluent Heun equation. We remark that the results of this paper apply to any physical problem governed by a Heun differential equation or a confluence of it. In particular, they can be easily adapted to the case of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes, cosmological perturbations [58]. It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^5] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to come back to this issue in a future publication.
@@ Acknowledgements
Appendix A Gauge/Gravity/CFT dictionary
In this Appendix, we display the dictionaries needed to translate the variables used in this paper to describe the confluent Heun equation to those used in gravity, gauge theory and in the CFT descriptions of the Heun equation. In all the cases, the confluent Heun can be written in the normal form
$\Psi^{\prime\prime}(z)+Q(z)\Psi=0$ (108)
with [^6] With respect to [30, 34] we find it convenient to redefine the coupling sending $x\to-x$ in $Q_{CFT}$.
$Q_{\rm CHE}$ $=$ ${-}\frac{x^{2}}{4z^{4}}{+}\frac{x{-}2x\kappa_{1}}{2z^{3}}{+}\frac{1{-}\left(\kappa_{2}{-}\kappa_{3}\right)^{2}}{4(z-1)z}{+}\frac{\left(2{-}\kappa_{2}{-}\kappa_{3}\right)\left(\kappa_{2}{+}\kappa_{3}\right)}{4(z{-}1)^{2}z}+\frac{u-{\textstyle\frac{1}{4}}{-}{\textstyle\frac{x}{2}}\left(\kappa_{2}{+}\kappa_{3}\right)}{(z{-}1)z^{2}}$
$Q_{\rm gauge}$ $=$ ${-}\frac{x^{2}}{4z^{4}}{+}\frac{xm_{3}}{z^{3}}{+}\frac{1{-}\left(m_{1}{-}m_{2}\right)^{2}}{4(z{-}1)z}+\frac{1{-}\left(m_{1}{+}m_{2}\right){}^{2}}{4(z-1)^{2}z}+\frac{u{-}{\textstyle\frac{1}{4}}{+}{\textstyle\frac{x}{2}}\left(m_{1}{+}m_{2}{-}1\right)}{(z{-}1)z^{2}}$
$Q_{\rm CFT}$ $=$ $-\frac{x^{2}}{4z^{4}}-\frac{cx}{z^{3}}+\frac{\frac{1}{4}-p_{0}^{2}}{(z-1)z}+\frac{\frac{1}{4}-k_{0}^{2}}{(z-1)^{2}z}+\frac{u-{\textstyle\frac{1}{4}}-x\left(\frac{1}{2}-k_{0}\right)}{(z-1)z^{2}}$ (109)
$Q_{\rm gravity}$ $=$ $\frac{4\omega^{2}M^{2}}{z^{4}}{-}\frac{8{\rm i}\omega M\left(1{+}{\rm i}\omega M\right)}{z^{3}}{-}\frac{3}{4(z-1)z}{+}\frac{16{\rm i}\omega M\left(1{-}{\rm i}\omega M\right){-}3}{4(z{-}1)^{2}z}{+}\frac{\ell(\ell{+}1){+}4{\rm i}\omega M\left(1{+}3{\rm i}\omega M\right)}{(z-1)z^{2}}$
Matching the five coefficients one finds the dictionaries
$\begin{array}[]{llll}{\rm Gauge}:&\kappa_{1}=\frac{1}{2}-m_{3}&\quad\kappa_{2}=\frac{1}{2}-m_{2}&\quad\kappa_{3}\to\frac{1}{2}-m_{1}\\ {\rm CFT}:&\kappa_{1}=\frac{1}{2}+c&\quad\kappa_{2}=\frac{1}{2}-k_{0}+p_{0}&\quad\kappa_{3}\to\frac{1}{2}-k_{0}-p_{0}\\ {\rm Gravity}:&\kappa_{1}=\frac{5}{2}+2{\rm i}\omega M&\quad\kappa_{2}=\frac{5}{2}-2{\rm i}\omega M&\quad\kappa_{3}=\frac{1}{2}-2{\rm i}\omega M\\ \end{array}$ (110)
with
$x=4{\rm i}\omega M\,,\,u\to(\ell+{\textstyle\frac{1}{2}})^{2}-4M^{2}\omega^{2}+10{\rm i}\omega M$ (111)

### v3

@@ 5.1 The hypergeometric representation
$G^{0}_{\alpha}({\textstyle\frac{x}{z}})g_{\alpha}(x)=G^{1}_{\alpha}(z)+\ldots$ (97)
@@
$H^{0}_{\alpha}(y)$ $=$ $y^{\frac{1}{2}-\alpha a-m_{3}}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}-\alpha a-m_{3},{\textstyle\frac{1}{2}}-\alpha a-m_{4};1-2\alpha a;y\right)$
$H^{1}_{\alpha}(z)$ $=$ $z^{-\frac{1}{2}+\alpha a+m_{3}}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}+\alpha a-m_{1},{\textstyle\frac{1}{2}}+\alpha a-m_{2};1+2\alpha a;z\right)$ (98)
@@
$P_{0}(y)$ $=$ $1+\sum_{i=1}^{k}\sum_{j=0}^{i-1}c_{ij}\,x^{i}y^{j-i}\quad,\quad\widehat{P}_{0}(y)=\sum_{i=1}^{k}\sum_{j=0}^{i}\widehat{c}_{ij}\,x^{i}y^{j-i}$
$P_{1}(z)$ $=$ $1+\sum_{j=1}^{k}\sum_{i=0}^{j-1}d_{ij}\,z^{i-j}x^{j}\quad,\quad\widehat{P}_{1}(z)=\sum_{j=1}^{k}\sum_{i=0}^{j}\widehat{d}_{ij}\,z^{i-j}x^{j}$
We notice that now only the non-confluent hypergeometric functions are involved. As before, plugging the ansätze (97) into (89) and setting to zero the coefficients of $H^{p}_{\alpha}$ ($p=0,1$) and its derivatives, one finds a linear system of algebraic equations for the coefficients $c_{ij}$, $\widehat{c}_{ij}$, $d_{ij}$, $\widehat{d}_{ij}$. For example, for $k=1$, writing $u=a^{2}+u_{1}x+\ldots$, one finds
$c_{10}$ $=$ $\frac{1}{2}(1-m_{1}-m_{2}-m_{3})-\frac{2m_{1}m_{2}m_{3}}{4a^{2}-1}\quad,\quad\widehat{c}_{11}=-\widehat{c}_{10}=\frac{1}{2}+\frac{2m_{1}m_{2}}{4a^{2}-1}$
$d_{01}$ $=$ $\frac{2\left(m_{3}-1\right)m_{3}m_{4}}{4a^{2}-1}-\frac{m_{4}}{2}\quad,\quad\widehat{d}_{11}=-\widehat{d}_{01}=\frac{2m_{3}m_{4}}{4a^{2}-1}+\frac{1}{2}$
$u_{1}$ $=$ $\frac{2m_{1}m_{2}m_{3}m_{4}}{4a^{2}-1}+\frac{1}{8}\left(4a^{2}-1\right)+\frac{1}{2}\left(1-\sum_{i=1}^{4}m_{i}+\sum_{i<j}^{4}m_{i}m_{j}\right)$ (100)
As in the confluent case, we denote by
the ratio of the two solutions. Explicit computations show that, as in the confluent case, the ratio of the two components can be written as
$\gamma(x)={g_{+}(x)\over g_{-}(x)}=x^{2a}e^{-\partial_{a}{\cal F}_{\rm inst}(a,x)}=x^{2a}\left(1+x\left(a-\frac{16am_{1}m_{2}m_{3}m_{4}}{\left(1-4a^{2}\right)^{2}}\right)+\ldots\right)$ (102)
with ${\cal F}_{\rm inst}$ still is given by (22). The function ${\cal F}_{\rm inst}$ corresponds now to the prepotential of a ${\cal N}=2$ supersymmetric $SU(2)$ gauge theory with four fundamental flavours.
5.2 Connection formulae
@@
$\widetilde{H}^{0}_{\alpha}(y)$ $=$ $(e^{{-}i\pi}y)^{{-}\frac{1{+}\alpha}{2}m_{34}}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}{-}a{-}m_{3}{+}{\textstyle\frac{1{+}\alpha}{2}}m_{34},{\textstyle\frac{1}{2}}{+}a{-}m_{3}{+}{\textstyle\frac{1{+}\alpha}{2}}m_{34};1{+}\alpha m_{34};{\textstyle\frac{1}{y}}\right)$ (103)
$\widetilde{H}^{1}_{\alpha}(z)$ $=$ $z^{a{+}m_{3}{-}\frac{1}{2}}(1{-}z)^{\frac{1+\alpha}{2}\left(m_{1}{+}m_{2}\right)}\,_{2}F_{1}\left({\textstyle\frac{1}{2}}{+}a{+}\alpha m_{1},{\textstyle\frac{1}{2}}{+}a{+}\alpha m_{2};1{+}\alpha\left(m_{1}+m_{2}\right);1{-}z\right)$
with $m_{34}=m_{3}-m_{4}$. We notice that both $\widetilde{H}^{0}_{\alpha}(y)$ and $\widetilde{H}^{1}_{\alpha}(z)$ are invariant under the sign flipping $a\to-a$, due to hypergeometric identities. The functions (103) are related to those introduced in (98) by the standard hypergeometric relations
$H^{0}_{\alpha}(y)=\sum_{\alpha^{\prime}=\pm}B_{\alpha\alpha^{\prime}}\widetilde{H}^{0}_{\alpha^{\prime}}(y)\quad,\quad H^{1}_{\alpha}(z)=\sum_{\alpha^{\prime}=\pm}F_{\alpha\alpha^{\prime}}\widetilde{H}^{1}_{\alpha^{\prime}}(z)$ (104)
@@
$B_{\alpha\alpha^{\prime}}$ $=$ $e^{-i\pi\left(\alpha a+m_{3}-\frac{1}{2}\right)}\frac{\Gamma(1-2\alpha a)\Gamma\left(\alpha^{\prime}\left(m_{4}-m_{3}\right)\right)}{\Gamma\left({\textstyle\frac{1}{2}}-\alpha a-\alpha^{\prime}m_{3}\right)\Gamma\left({\textstyle\frac{1}{2}}-\alpha a+\alpha^{\prime}m_{4}\right)}$
$F_{\alpha\alpha^{\prime}}$ $=$ $\frac{\Gamma(1+2\alpha a)\Gamma\left(-\alpha^{\prime}\left(m_{1}+m_{2}\right)\right)}{\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{1}\right)\Gamma\left({\textstyle\frac{1}{2}}+\alpha a-\alpha^{\prime}m_{2}\right)}$ (105)
Plugging (104) into (5.1) one finds the connection formulae
@@
$\widetilde{G}^{0}_{\alpha}(y)$ $=$ $P_{0}(y)\widetilde{H}^{0}_{\alpha}(y){+}\widehat{P}_{0}({\textstyle\frac{x}{z}})y\widetilde{H}^{0}_{\alpha}{}^{\prime}(y)$
@@
6 Summary and conclusions
In this paper, we derive a hypergeometric representation of the confluent Heun functions $G_{0}(y)$ or $G_{1}(z)$ that resums the entire $z$ or $y$ dependence of the functions respectively, at any given order in the expansion for small $x$. In the context of gravity, these representations provide exact formulae for the wave form at given order in the PM or PN approximations respectively. The results apply to a general scattering process (along an open or closed trajectory) in the Schwarzschild geometry, where the expansion parameters $z=2M/r$ and $y=2{\rm i}\omega r$ are independent from each other. In a future publication we are going to compare this result with those coming from amplitude computations which, at a given loop order are exact in the velocities. We build four equivalent basis of solutions, $\widetilde{G}^{1}_{\alpha}$, $G^{1}_{\alpha}(z)$, $G^{0}_{\alpha}(z)$ and $\widetilde{G}^{0}_{\alpha}$ that describe with good accuracy different regions of the space time, and are connected to each other by Heun connection matrices. The final picture is summarized in the following scheme
@@ 6 Summary and conclusions
$\begin{array}[]{lll}{\rm Near~{}horizon}&{\rm Near~{}zone}&{\rm Wave~{}zone}\\ z\approx 1&x\ll z\ll 1&z\ll x\ll 1\\ F_{\alpha\alpha^{\prime}}\widetilde{G}^{1}_{\alpha^{\prime}}&G^{1}_{\alpha}=g_{\alpha}G^{0}_{\alpha}&g_{\alpha}B_{\alpha\alpha^{\prime}}\widetilde{G}^{0}_{\alpha^{\prime}}\end{array}$ (112)
@@
where $B_{\alpha\alpha^{\prime}}$ and $F_{\alpha\alpha^{\prime}}$ are the standard hypergeometric connection matrices relating $z$ to $1/z$ or $1-z$ respectively. The whole non-triviality of the Heun connection formula is codified in the dependence of these matrices on the characteristic function $a(\ell)$. This simple result is a direct consequence of our ansatz for the solutio of the differential equation in terms of a single hypergeometric function and its derivative.
@@
Specifying to circular orbits, we derive the wave form and energy flux at order 30PN. Given that this result is too large to appear in print we also publish ancillary mathematica files with the data needed to perform the computation. When compared against the PN expansions that one can find in literature Blanchet:2013haa ; Fujita:2012cm , these formulae resum the infinite tower of logarithmically divergent terms (tail and tail of tails), responsible for the slow convergence of the PN series, into exponentials. They also summarise in a compact form the dependence on $\ell$ of the waveforms, all the so called tail and tail of tail terms of the multipole expansion and have no logarithmic terms (responsible for the slow convergence of the PN series). The results for the waveform have been compared against those obtained using the numerical package BHToolkit, finding an agreement from 9 to 12 digits in the whole range of parameters allowed for bounded orbits.
@@
Similar results can be also obtained for solutions of the non-confluent Heun equation and can therefore be applied to other physical problem governed by a Heun differential equation or a confluence of it. In particular, they can be easily adapted to the case of collisions of spinning BHs (Kerr geometry) or orbital motions in the backgrounds of topological stars, D-branes, cosmological perturbations Bianchi:2021xpr ; Bianchi:2021mft ; DiRusso:2024hmd ; Bianchi:2024vmi ; Bianchi:2024mlq ; Bianchi:2024rod . It would also be very interesting to apply our results to the generation of templates of waveforms for GWs data analysis [^8] We thank A.Nagar and M.Panzeri for useful discussions on this issue.. We hope to also come back to this issue in a future publication.